OpenStudy (anonymous):

what is coordinates of the midpoints of (0,0) (8,-4) and (4,-8) (8,-4)

OpenStudy (anonymous):

for mid-point use (x1+x2)/2 ,(y1+y2)/2 hence for first pair it is (0+8)/2 ,(0-4)/2 =4,-2

OpenStudy (lukebluefive):

Add each point pairs' respective coordinates to each other, then divide by 2. This gives the following coordinates: Point 1: (4, -2) Point 2: (6, -6)

OpenStudy (anonymous):

for the second pair similrly u gt (4+8)/2,(-8-4)/2 =6,-6

OpenStudy (anonymous):

the qustion of this hole things is suppose a triangular component in CAD display of different tower has vertices A)(0,0) B)(4,-8) C)(8,-4) so there for what are the coordinates of the midpoint of (0,0) (8,-4) and (4,-8) (8,-4)

OpenStudy (lukebluefive):

Wouldn't you need the midpoint for (0, 0) and (4, -8) as well?

OpenStudy (anonymous):

yeah but how do u do all this

OpenStudy (lukebluefive):

The midpoint of line AB is found via the following: \[\frac{(0+4)}{2}=2, \frac{(0-8)}{2}=-4\] Basically, since the coordinate points are written in the form (x, y), add both of the x-coordinates together and divide by 2, then add both of the y-coordinates together and divide by 2 again. The first gives you the x-coordinate of the midpoint, while the second gives you the y-coordinate.

OpenStudy (anonymous):

thanks you